# Benchmarking Post-Training Quantization in LLMs: Comprehensive Taxonomy, Unified Evaluation, and Comparative Analysis

### v1

@@ 3.2 Cross Bitwidth and Training Level Robustness
Extremely Low-bit PTQ Performance Varies with Training Level of LLMs.
As the technical report (Dubey et al., 2024), the training dataset of LLaMA3/3.1 is several times larger than that of LLaMA/LLaMA2. The more extensive training leads to greater information loss when quantizing LLaMA3/3.1, especially for extremely low-bit PTQ, i.e., 2-bit. Our experimental results in Table 1 and Table 2 substantiate this inference and align with the recently proposed quantization scaling law (Kumar et al., 2024; Ouyang et al., 2024). However, previous research ignores to explore an effective PTQ pipeline for knowledge intensive LLMs. In this benchmark, we give a comprehensive insight through extensive experiments on undertrained LLMs (Table 1) and fully trained LLMs (Table 2) (Ouyang et al., 2024). According to the results, we observe that the performance of each PTQ strategy varies significantly with changes in the training level of LLMs at 2-bit, as summarized in three dimensions below:
@@
Salience-based methods collapse on All LLMs: As listed in Table 1, AWQ exhibits extremely poor performance at 2-bit across all LLMs, completely losing any language capabilities, e.g., on LLaMA-65B AWQ only delivers 7.4e4/22.86%..
@@
Optimization-based methods collapse on fully trained LLMs: Although OmniQuant performs well on LLaMA/LLaMA2, its performance drastically declines on more extensively trained LLaMA3/3.1, particularly at 3-bit where other baselines still maintain decent performance, OmniQuant has already collapsed. For instance, on 3-bit LLaMA3-70B, OmniQuant achieves 7.4e4/22.22%, whereas GPTQ is 7.16/40.51%.
@@
Rotation-based and Compensation-based methods demonstrate low-bit and cross training level robustness: It is evident that GPTQ and QuIP consistently achieve satisfactory results across any model and bitwidth. For example, GPTQ scores 19.50/29.55 while AWQ scores 1.1e5/22.61 on 2-bit LLaMA2-13B, and on 2-bit LLaMA3.1-70B, QuIP is 42.58/25.58, whereas OmniQuant achieves 1.6e5/21.85. Specifically, QuIP is more suitable for undertrained LLMs, while GPTQ demonstrates stronger robustness for 2-bit quantization of fully trained LLMs. For instance, QuIP (7.70/48.47) outperforms GPTQ (9.00/43.44) on LLaMA2-70B while on LLaMA3-70B GPTQ (23.43/29.9%) surpasses QuIP (53.18/24.60). Overall, the observed phenomenon suggests that both compensation and rotation methods are applicable for low-bit PTQ, with rotation is better suited for undertrained LLMs and compensation is more advantageous for fully trained LLMs.
3.3 Cross-Architecture Robustness
@@
LLMs constructed by stacking traditional transformer modules suffer from drawbacks such as high computational costs and poor long-sequence modeling capability. Consequently, various novel architectures for LLMs have emerged in recent years, such as MoE (Fedus et al., 2022) and Mamba (Gu and Dao, 2023). However, changes in model architecture may lead to potential issues of algorithm compatibility or performance degradation. Therefore, in this benchmark we further evaluate the four PTQ strategies on novel-architecture LLMs to analyze their cross-architecture robustness.
Methods W4 W3 W2
Mamba- 1.4B 1.4B 2.8B 1.4B 2.8B
AWQ - - - - -
GPTQ 12.74/43.12 15.64/40.18 13.35/43.84 882.35/24.53 696.21/24.06
QuIP 14.53/41.63 17.64/38.79 14.15/41.97 287.48/25.24 119.95/28.23
OmniQ 4.2e4/25.66 1.6e3/27.36 36.92/34.93 2.9e4/23.05 1.2e4/23.23
Table 3: The average perplexity($\downarrow$)/accuracy($\uparrow$) comparison among different bitwidths and model sizes of Mamba.
@@
Table 4: The average perplexity($\downarrow$)/accuracy($\uparrow$) comparison results among different bitwidths of MoE LLMs.
The evaluation results on Mamba 1.4B and 2.8B (Gu and Dao, 2023) are presented in Table 3. Mixtral 8$\times$7B (Jiang et al., 2024) and DeepSeekMoE-16B (Dai et al., 2024) are chosen for benchmarking MoE LLMs, where the results are shown in Table 4. For more details data please refer to Appendix B.2.
@@
As shown in Table 3, OmniQuant completely collapses on Mamba LLMs. Meanwhile, on Mixtral 8$\times$7B and DeepSeekMoE-16B, its performance is also unsatisfactory, particularly at 2-bit. The aforementioned two phenomena demonstrate that AWQ and OmniQuant exhibit poor cross-architecture generalization capabilities.
@@
The results in Table 3 and Table 4 indicate that GPTQ and QuIP achieve superior performance across various bitwidths for both MoE LLMs and Mamba. Taking the performance on LLaMA into consideration, it is evident that these two PTQ strategies maintain stability across different model-architecture. Interestingly, their robustness exhibit distinct nature. Specifically, in most cases, GPTQ performs better at higher bit, while QuIP is more suitable for 2-bit quantization. For example, GPTQ achieves 11.30/45.69% on Mamba-2.8B while QuIP is 12.30/45.11% at W4, but at W2 QuIP outperforms GPTQ with 119.95/28.23%. And on 2-bit DeepSeekMoE-16B, QuIP shows significant advantages with 23.29/32.79% compared with GPTQ. In addition, we can notice that QuIP exhibits unstable performance on MoE LLMs, especially on Mixtral 8$\times$7B where GPTQ significantly outperforms QuIP across all quantization bitwidths. For instance, GPTQ and QuIP exhibit 22.92/28.19% and 155.14/23.52% on Mixtral at W2, respectively. The observed phenomenon suggests that compensation-based methods represented by GPTQ is the most cross-architecture robust strategy but rotation-based strategy excels particularly at 2-bit PTQ in most cases.
Methods W4 W3 W2
VILA1.5- 8B 8B 13B 8B 13B
AWQ 61.23 52.77 68.36 16.90 8.65
GPTQ 62.55 52.77 67.72 18.70 42.59
QuIP 58.32 54.29 67.30 21.50 59.42
OmniQ 58.05 41.70 66.06 17.54 9.42
LLaVA1.5- 7B 7B 13B 7B 13B
AWQ 59.94 58.68 62.24 8.56 8.74
GPTQ 59.67 57.18 61.79 14.40 34.87
QuIP 59.24 53.89 62.12 21.99 42.24
OmniQ 59.21 54.54 60.33 9.67 8.78
Table 5: The average accuracy (%, $\uparrow$) comparison among different bitwidths and model sizes of MLLMs.
@@ 3.4 Cross-Modality Robustness
Quantization may undermine the inherent cross-modal alignment capabilities, leading to degraded performance on multimodal tasks. In this benchmark, we further evaluate the average accuracy of the four PTQ strategies in visual-language reasoning tasks on VILA (Lin et al., 2023) and LLaVA (Liu et al., 2024a) to explore cross-modality robustness, and the results are presented in Table 5. For more details please refer to Appendix B.2.
@@
As shown in Table 5, the phenomenon at 2-bit is consistent with that in LLMs, where only models quantized by GPTQ and QuIP exhibit effective reasoning capabilities while AWQ and OmniQuant completely collapse. For instance, on VILA1.5-8B AWQ shows 7.69% and QuIP remains 54.18%. The experimental observations suggest that at higher bit, any strategy exhibits commendable cross-modal robustness, whereas at extremely low-bit, rotation-based and compensation-based strategies emerge as the sole viable alternatives.
@@ 3.5 Trade-off among Bitwidth, Model Size and Performance
When deploying quantized LLMs in real system, people often grapple with the decision regarding the size and the bitwidth of the model to deploy. For instance, a straightforward question arises: Which one is better, a higher-bit smaller model or a lower-bit larger model? In order to clarify this confusion, in this benchmark, we comprehensively explore the trade-off between model size and quantization bitwidth from the performance perspective. The experimental results in previous tables indicate that the perplexity of text generation exhibits a strictly positive correlation with reasoning ability reflected in accuracy. As perplexity exhibits a much larger distribution range, we choose accuracy as the evaluation metric and then create intuitive visualizations to explore the trade-off (see Figure 2).
@@
The scaling law claims that, within the same LLM family, larger models generally exhibit superior performance. However, Figure 2 indicates that even the largest models, like LLaMA-65B, when quantized to extremely low-bit (2-bit), demonstrate inferior performance compared to the smallest models operating at 4-bit, such as LLaMA-7B. This observation remains valid across any LLM family, any architecture and any PTQ strategy as the results in the tables above. For instance, LLaMA2-70B quantized to 2-bit using OmniQuant achieves 10.36/38.46%, which is worse than LLaMA2-7B at 4-bit with 6.55/51.11%; Mamba-2.8B at 2-bit is also worse than Mamba-1.4B at 4-bit quantized by any baseline. This fresh finding suggests that for the present moment, deploying smaller models at higher bitwidths appears to be the optimal choice. Additionally, specific research on extremely low-bit PTQ for ultra-large models seems essential. Most previous PTQ methods usually validated their performance across various model sizes, which ignored the fact that pushing ultra-large models to extremely low-bit does not surpass the performance of higher-bit smaller models so as to cause the unnecessary expenditure of computational resources. This highlights the need for the development of specialized designs for extremely low-bit PTQ on ultra-large models to ensure their performance exceeds that of higher-bit small models.
@@
4 Compensation-based PTQ: A Unified Robust Foundational Strategy
In Section 3, we have provided several recommendations for the selection of foundational PTQ strategy based on different requirements. But when developing novel PTQ algorithm, researchers usually strive for an ideal form, one where the designed method can be effectively generalized to any quantization scenario. To address this issue, in this section we further explore this valuable requirement to improve our benchmark.
Based on the results in Section 3, it is evident that rotation-based and compensation-based strategies exhibit better generalization ability. Subsequently, we broadened our focus to more advanced algorithms within compensation-based and rotation-based strategies, i.e., VPTQ and QuaRot. However, when validating on Mamba and MoE LLMs, we discovered that QuaRot (Ashkboos et al., 2024), the state-of-the-art rotation-based method, could not be applied to Mamba and MoE LLMs, because QuaRot introduces additional rotation matrices that need to be integrated into the nearby linear layers. However, due to changes in the model architecture, this integration process cannot be achieved. In contrast, any compensation-based method remains unaffected , as this strategy only relies on the weights themselves to complete updates and error compensation.
Subsequently, we combine GPTQ with other strategies and compare them with more advanced algorithms in other strategies to highlight the superiority in challenging scenarios including extremely low-bit (2-bit), high training level (LLaMA3.1-8B), and cross-architecture (Mixtral 8$\times$7B) settings. We select naive GPTQ, OWQ (SOTA salience-based method) (Lee et al., 2024), LRQuant (SOTA optimization-based method) (Zhao et al., 2024), QuaRot+GPTQ (rotation+compensation), and PBLLM (salience+compensation) (Shang et al., 2023). Table 6 reveals the PBLLM and QuaRot+GPTQ exhibit consistently high stability, outperforming more advanced salience-based and optimization-based methods as well as naive GPTQ. Our exploration claims that compensation-based PTQ is the most unified robust strategy, and its combination with other strategies can significantly elevate the performance ceiling.
Model Methods PPL $\downarrow$ Acc $\uparrow$
LLaMA3.1-8B GPTQ 510.33 23.20
OWQ 305.77 24.33
LRQuant 193.83 26.05
PBLLM 84.45 33.27
QuaRot+GPTQ 33.33 35.28
Mixtral 8$\times$7B GPTQ 22.92 28.19
OWQ 35.60 32.89
LRQuant 41.81 29.77
QuIP+GPTQ 9.59 45.26
PBLLM 7.60 56.28
Table 6: The average perplexity($\downarrow$)/accuracy($\uparrow$) comparison among different PTQ strategy combination.
@@ 5 Conclusion
In this paper, we introduce a novel comprehensive benchmark for PTQ methods in LLMs. Specifically, our benchmark first provide a comprehensive taxonomy for existing mainstream PTQ methods. Then with the baselines of each PTQ strategy, we conduct extensive experiments on LLMs with various training levels, model sizes, architectures and modalities. According to the evaluation results, we summarize the characteristic of each PTQ strategy and discover the modelsize-bitwidth trade-off. At last, with the comparative analysis, our benchmark offers valuable guidelines for quantized LLMs deployment and future PTQ algorithms development.
Limitation
Our benchmark addresses the shortcomings of existing PTQ surveys and provides guidance for future research. However, due to the rapid development and vast number of PTQ algorithms, it is impractical for us to conduct experiments using every approach across various models and scenarios to generate more comprehensive results for summarization. Additionally, because of differences in experimental environments, our experimental results may deviate from those originally reported. Nevertheless, to ensure fairness, all our experimental settings are identical. Lastly, our benchmark solely covers weight-only quantization, as the characteristics and strategies for weight-activation quantization are entirely different, and including them would significantly lengthen this paper. In the future, we will propose a specific benchmark for weight-activation quantization to fill the gap in quantization community.
Ethics Statement
This paper presents in-depth insights and recommendations associated with Large Language Models (LLMs) quantization, with the overarching goal of facilitating the widespread adoption and application of LLMs. In the current landscape, ethical concerns tied to LLMs, including the presence of hidden biases encoded in the models, are garnering heightened attention. Following our investigation, we assert that our proposed method does not further amplify the biases and contravene any ethical standards.

### v4

@@ 3.2 Cross-Bitwidth Robustness
Extremely Low-bit PTQ Performance Varies
According to the results in Table 1 and Table 2, we observe that the performance of each PTQ strategy varies significantly at 2-bit, as summarized in three dimensions below:
@@
- Salience-based methods collapse on All LLMs: As listed in Table 1 and Table 2, AWQ exhibits extremely poor performance at 2-bit across all LLMs, completely losing any language capabilities, e.g., on LLaMA-65B AWQ only delivers 7.4e4/22.86%.
@@
- Optimization-based methods collapse on LLaMA-3/3.1: Although OmniQuant performs well on LLaMA-1/2, its performance drastically declines on LLaMA-3/3.1, particularly at 3-bit where other baselines still maintain decent performance, OmniQuant has already collapsed. For instance, on 3-bit LLaMA3-70B, OmniQuant achieves 7.4e4/22.22%, whereas GPTQ is 7.16/40.51%.
@@
- Rotation-based and Compensation-based methods demonstrate diverse low-bit robustness on LLaMA-1/2 and LLaMA-3/3.1: It is evident that GPTQ and QuIP consistently achieve satisfactory results across any model and bitwidth. For example, GPTQ scores 19.50/29.55% while AWQ scores 1.1e5/22.61% on 2-bit LLaMA2-13B, and on 2-bit LLaMA3.1-70B, QuIP is 42.58/25.58%, whereas OmniQuant achieves 1.6e5/21.85%. Specifically, QuIP is more suitable for LLaMA-1/2, while GPTQ demonstrates stronger robustness for 2-bit quantization of LLaMA-3/3.1. For instance, QuIP (7.70/48.47%) outperforms GPTQ (9.00/43.44%) on LLaMA2-70B while on LLaMA3-70B GPTQ (23.43/29.9%) surpasses QuIP (53.18/24.60%). To elucidate this multifaceted phenomenon, we provide the following analysis: As the technical report [16], the training dataset of LLaMA-3/3.1 is several times larger than that of LLaMA-1/2. The more extensive training leads to greater information loss when quantizing LLaMA-3/3.1, especially for extremely low-bit PTQ (2-bit). Our experimental results in Table 1 and Table 2 substantiate this inference and align with the recently proposed quantization scaling law [33, 53].
Overall, the observation in low-bit evaluation suggests that both compensation and rotation methods are applicable for low-bit PTQ, with rotation is better suited for undertrained LLMs (LLaMA-1/2) and compensation is more advantageous for fully trained LLMs (LLaMA-3/3.1).
3.3 Cross-Structure Robustness
@@
LLMs constructed by stacking traditional transformer modules suffer from high computational costs and poor long-sequence modeling capability. Consequently, various novel structures for LLMs have emerged recently, such as MoE [17] and Mamba [24]. However, changes in model structure may lead to potential issues of algorithm compatibility or performance degradation. Therefore, in PTQ-Bench we further evaluate the four PTQ strategies on novel-structure LLMs to analyze their cross-structure robustness. Specifically, we select Mamba-1.4B/2.8B as the representives of Mamba families, while Mixtral 8$\times$7B [32] and DeepSeekMoE-16B [12] are chosen for benchmarking MoE LLMs. The results are shown in Table 3. For more details data please refer to Appendix B.2.
[…]
@@
Mamba-1.4B AWQ - - -
GPTQ 12.74/43.12 15.64/40.18 882.35/24.53
QuIP 14.53/41.63 17.64/38.79 287.48/25.24
OmniQ 4.2e4/25.66 1.6e3/27.36 2.9e4/23.05
Mamba-2.8B AWQ - - -
GPTQ 11.30/45.69 13.35/43.84 696.21/24.06
QuIP 12.30/45.11 14.15/41.97 119.95/28.23
OmniQ 20.16/38.28 36.92/34.93 1.2e4/23.23
@@
As shown in Table 3, OmniQuant completely collapses on Mamba LLMs. Meanwhile, on Mixtral 8$\times$7B and DeepSeekMoE-16B, its performance is also unsatisfactory, particularly at 2-bit. The aforementioned two phenomena demonstrate that AWQ and OmniQuant exhibit poor cross-structure generalization capabilities.
@@
The results in Table 3 indicate that GPTQ and QuIP achieve superior performance across bitwidths for both MoE LLMs and Mamba. Taking the performance on LLaMA into consideration, it is evident that these two PTQ strategies maintain stability across different model-structure. Interestingly, their robustness exhibit distinct nature. Specifically, in most cases, GPTQ performs better at higher bit, while QuIP is more suitable for 2-bit quantization. For example, GPTQ achieves 11.30/45.69% on Mamba-2.8B (W4) while QuIP is 12.30/45.11%, but at W2 QuIP outperforms GPTQ with 119.95/28.23%. And on 2-bit DeepSeekMoE-16B, QuIP shows significant advantages with 23.29/32.79% compared with GPTQ. In addition, we can notice that QuIP exhibits unstable performance on MoE LLMs, especially on Mixtral 8$\times$7B where GPTQ significantly outperforms QuIP across all quantization bitwidths. For instance, GPTQ and QuIP exhibit 22.92/28.19% and 155.14/23.52% on Mixtral at W2, respectively. The observed phenomenon suggests that compensation-based is the most cross-structure robust strategy but rotation-based strategy excels particularly at 2-bit PTQ in most cases.
@@ 3.4 Cross-Modality Robustness
Quantization may undermine the inherent cross-modal alignment capabilities, leading to degraded performance on multimodal tasks. In PTQ-Bench, we further evaluate the four PTQ strategies in visual-language reasoning tasks on VILA [41] and LLaVA [43] to explore cross-modality robustness, and the results are presented in Table 4. For more details please refer to Appendix B.2.
@@
As shown in Table 4, the phenomenon at 2-bit is consistent with that in LLMs, where only models quantized by GPTQ and QuIP exhibit effective reasoning capabilities while AWQ and OmniQuant completely collapse. For instance, on VILA1.5-13B AWQ shows 7.69% and QuIP remains 54.18%. The experimental observations suggest that at higher bit, any strategy exhibits commendable cross-modal robustness, whereas at extremely low-bit, rotation-based and compensation-based strategies emerge as the sole viable alternatives.
After all evaluation, we theoretically analyze the reason why certain methods fail or work better in specific scenarios for providing deeper insights into these PTQ strategies, please refer to Appendix C.
[…]
Methods W4 W3 W2
VILA1.5- 7B 7B 13B 7B 13B
AWQ 68.55 66.30 69.91 7.56 7.69
GPTQ 68.59 65.95 69.33 19.30 45.93
QuIP 63.52 64.46 68.96 29.42 54.17
OmniQ 67.71 65.37 67.68 19.99 8.39
LLaVA1.5- 7B 7B 13B 7B 13B
AWQ 61.97 60.76 64.08 7.61 7.77
GPTQ 61.74 59.36 63.75 17.45 38.46
QuIP 61.32 56.18 64.02 24.80 45.59
OmniQ 61.27 56.94 62.38 8.61 7.80
4 Beyond Benchmarking: Extended Insights and Implications
In addition to the comprehensive evaluation for the categorized PTQ strategies and recommendations in different scenarios in Section 3, in this section we present two innovative and critical implications deriving from PTQ-Bench: a) Modelsize-bitwidth tradeoff; b) compensation is the unified robust foundational strategy.
@@ 4.1 Trade-off among Bitwidth, Model Size and Performance
When deploying quantized LLMs in real system, people often grapple with the decision regarding the size and the bitwidth of the model to deploy. For instance, a straightforward question arises: Which one is better, a higher-bit smaller model or a lower-bit larger model? In order to clarify this confusion, we comprehensively explore the trade-off between model size and quantization bitwidth from the performance perspective. The experimental results in previous tables indicate that the perplexity of text generation exhibits a strictly positive correlation with reasoning ability reflected in accuracy. As perplexity exhibits a much larger distribution range, we choose accuracy as the evaluation metric and then create intuitive visualizations to explore the trade-off (see Figure 2).
@@
The scaling law claims that, within the same LLM family, larger models generally exhibit superior performance. However, Figure 2 indicates that even the largest models, like LLaMA-65B, when quantized to extremely low-bit (2-bit), demonstrate inferior performance compared to the smallest models operating at 4-bit, such as LLaMA-7B. This observation remains valid across any LLM family, any structure and any PTQ strategy as the results in the tables above. For instance, LLaMA2-70B quantized to 2-bit using OmniQuant achieves 10.36/38.46%, which is worse than LLaMA2-7B at 4-bit with 6.55/51.11%; Mamba-2.8B at 2-bit is also worse than Mamba-1.4B at 4-bit quantized by any baseline. This fresh finding suggests that for the present moment, deploying smaller models at higher bitwidths appears to be the optimal choice. Additionally, specific research on extremely low-bit PTQ for ultra-large models seems essential. Most previous PTQ methods usually validated their performance across various model sizes, which ignored the fact that pushing ultra-large models to extremely low-bit does not surpass the performance of higher-bit smaller models so as to cause the unnecessary expenditure of computational resources. This highlights the need for the development of specialized designs for extremely low-bit PTQ on ultra-large models to ensure their performance exceeds that of higher-bit small models.
@@
[…]
LLaMA GPTQ Compen. 510.33 23.20
3.1-8B OWQ Salien. 305.77 24.33
LRQuant Optim. 193.83 26.05
PBLLM Compen.+Salien. 84.45 33.27
QuaRot+GPTQ Compen.+Rot. 33.33 35.28
Mixtral GPTQ Compen. 22.92 28.19
8$\times$7B OWQ Salien. 35.60 32.89
LRQuant Optim. 41.81 29.77
QuIP+GPTQ Compen.+Rot. 9.59 45.26
PBLLM Compen.+Salien. 7.60 56.28
4.2 Compensation-based PTQ: A Unified Robust Foundational Strategy
In Section 3, we have provided several recommendations for the selection of foundational PTQ strategy based on different requirements. But when developing novel PTQ algorithm, researchers usually strive for an ideal form, one where the designed method can be effectively generalized to any quantization scenario. To address this issue, in this part we further explore this valuable requirement to improve our benchmark.
Based on the results in Section 3, it is evident that rotation-based and compensation-based strategies exhibit better generalization ability. Subsequently, we broaden our focus to more advanced algorithms within compensation-based and rotation-based strategies, i.e., VPTQ and QuaRot. However, when validating on Mamba and MoE LLMs, we discovered that QuaRot [4], the state-of-the-art rotation-based method, could not be applied to Mamba and MoE LLMs, because QuaRot introduces additional rotation matrices that need to be integrated into the nearby linear layers. However, due to changes in the model structure, this integration process cannot be achieved. In contrast, any compensation-based method remains unaffected, as this strategy only relies on the weights themselves to complete error compensation.
Subsequently, we combine GPTQ with other strategies and compare them with more advanced algorithms in other strategies to highlight the superiority in challenging scenarios including extremely low-bit (2-bit), high training level (LLaMA3.1-8B), and cross-structure (Mixtral 8$\times$7B) settings. We select naive GPTQ, OWQ (SOTA salience-based method) [37], LRQuant (SOTA optimization-based method) [79], QuaRot+GPTQ (rotation+compensation), and PBLLM (salience+compensation) [59]. Table 5 reveals the PBLLM and QuaRot+GPTQ exhibit consistently high stability, outperforming more advanced salience-based and optimization-based methods as well as naive GPTQ. Our exploration claims that compensation-based PTQ is the most unified robust strategy, and its combination with other strategies can significantly elevate the performance ceiling.
@@ 5 Conclusion
Existing PTQ research lacks in-depth analysis of the superiority and applicable scenarios of various PTQ frameworks, presenting a challenge for future researchers in selecting foundational frameworks that align with their specific requirements for development. To address this issue, in this paper we first propose a comprehensive taxonomy for existing baselines. Then we establish a novel benchmark for LLMs PTQ, named PTQ-Bench, to help select foundational strategies by evaluating the cross-bitwidth, cross-structure and cross-modality robustness of the categorized strategies on a wide range of LLMs and metrics. Based on the results, we provide a series of analysis and recommendations. For example, rotation-based method shows excellent low-bit robustness. In addition, we also discover two critical insights deriving from PTQ-Bench including modelsize-bitwidth tradeoff and the most robustness strategy combination across various scenarios. We believe our PTQ-Bench will provide valuable guidelines for future development on PTQ methods.